OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 2x - 3y = -18. Answers: a: x intercept = (9, 0); y intercept = (0, -6) b: x intercept = (6, 0); y intercept = (0, -9) c: x intercept = (-9, 0); y intercept = (0, 6) d: x intercept = (-6, 0); y intercept = (0, 9)

OpenStudy (anonymous):

The X-Intercept is where the linear equation passes through the X-axis. Touching the x-axis means that the y-value needs to be zero. So, to solve for the x-intercept, we substitute (0) in for y, and solve. \[2x-3y=-18\] \[2x-3(0)=-18\] \[2x-0=-18\] \[2x=-18\] \[x=-9\] So for the y-value of 0, the x-value is -9, so the x-intercept is the ordered pair (-9,0) The Y-Intercept is where the linear equation passes through the Y-axis. Passing through the y-axis means that the x-value needs to be zero. So, to solve for the y-intercept, we substitute (0) in for x, and solve. \[2x-3y=-18\] \[2(0)-3y=-18\] \[-3y=-18\] \[y=6\] So for the x-value of 0, the y-value is 6, so the y-intercept is the ordered pair (0,6) In y our multiple choice, you see this represented as choice C